# Machine-assisted classification of potential biosignatures in earth-like exoplanets using low signal-to-noise ratio transmission spectra

### v1

@@ 6.1.3 BC interference analysis
We can use similar methods as before to evaluate the performance of the algorithm in classifying non-bio planets with H2O, as interesting cases. This is, to measure the tendency to misclassify planets due to the presence of water. As we observe in Figure 2, the water bands in the spectral range we are considering here overlap most of the ozone and some of the methane bands. As a result, when we have a large water mixing ratio and a low SNR, a non-bio planet can be tagged as interesting.
In Figure 5 we show the value of TNR (which is the best-suited metric for this case) as a function of water mixing ratio. As expected, the TNR is reduced, i.e. the algorithm is not tagging as not-interesting planets without biosignatures (TN), when the mixing ratio of water is increased. When the SNR is increased, the performance of the algorithm improves in the sense of becoming less sensitive to this water inteference.
@@
Figure 5: TNR in BC across H2O mixing ratios evaluated at representative values of SNR. The shaded area below 0.6 in the graphs indicates suboptimal classification performance.
@@
We also analyse the performance of the algorithm with planets having two biosignatures. For that purpose, we picked samples of spectra having pairs of mixing ratios such as (CH4, O3) : $(10^{-10},10^{-10})$, $(10^{-10},10^{-9}),\ldots$, $(10^{-9},10^{-10})$, $(10^{-9},10^{-9}),\ldots$. For each sample, we compute the Recall. In the columns of Figure 6 we show the result as heat map of this metric as a function of mixing ratio combinations. We call this Recall maps.
@@
Figure 6: Grid of Recall maps (see text) ssfor BC at two repersentative SNR levels (2 and 10). Recall was calculated for different pairs of molecules across varying mixing ratios
@@
As we observe even at low mixing ratios of one of the biosignatures, the Recall can be high if the mixing ratio of the other one is high enough. For example, at SNR = 2 (first row in Figure 6) and at a very low concentration of ozone, e.g. 10-9 in the first panel, the presence of abundant methane, e.g. 10-3 will make that the algorithm classify as interesting most planets having both biosignatures. Our interpretation of this result is that the algorithm seems to learn which are the spectral bands of each biosignature independently and it is able to know that the planet is interesting by identifying any of them even if the other is missing.
@@
Additionally, as SNR increases, the algorithm becomes more conservative (lower values of Recall) when both mixing ratios are low but achieves better performance (higher values of Recall) at higher mixing ratios.
@@
We repeated the same procedure as before but now using combinations of one of the biosignatures and water. The resulting Recall maps are shown in the second and third columns of Figure 6. Although water tends to interfere in the detection of biosignatures, the fact that the frontier between low and high Recall areas in the Recall maps is almost horizontal, means that what determines the performance of the algorithm is mainly the mixing ratio of the biosignature. In other words, H2O only affects negatively the Recall in situations where the mixing ratio of the biosignature is really low.
6.2 Multilabel Classificator
@@
In our previous experiment we binary classified the planets as interesting and not-interesting. We aim now to improve the capacity of the method by introducing multiple labels, intended to identify individual biosignatures and other interesting molecules.
For that purpose, we specifically define three labels: methane, ozone and water, each of them having two possible values: 0 (the molecule is absent) and 1 (the molecule is present).
@@
Focusing on the detection of each molecule individually is valuable because, as discussed in section 4, different biosignatures have varying degrees of robustness and implications. By independently classifying each biosignature, we can better understand the presence and relevance of each molecule, thereby improving our ability to assess the potential habitability of exoplanets.
@@
The MC random forest differs in its hyperparameters from the BC algorithm described in subsection 6.1 by having a lower number of estimators (100) and a defined max depth of 200. The rest of the hyperparameters remain the same.
@@
Table 5: Training data distribution for the multilabel classificator (MC). This table has the same structure as Table 4.
SNR 000 100 010 001 110 101 011 111
1 10000 250 250 250 50 50 50 5
3 10000 250 250 250 50 50 50 5
6 5000 250 250 250 50 50 50 5
10 5000 250 250 250 50 50 50 5
20 5000 250 250 250 50 50 50 5
0 10000 250 250 250 50 50 50 5
After training, the classification thresholds for CH4, O3, and H2O were set at $0.49$, $0.45$, and $0.50$, respectively.
The model was tested by generating 150 000 noisy spectra per SNR, including 30 000 spectra for the scenario without interesting molecules, and another 30 000 for the cases where all three molecules are present. For the remaining combinations, we generate 15 000 spectra per combination for a total of 90 000. This distribution ensures a balanced test dataset, i.e. for each molecule, there is an equal number of spectra with and without it.
@@ 6.2.1 MC multilabel metrics
The algorithm improves more rapidly in its ability to avoid false positives as measured by TNR and Precision when the SNR increases. This means that, on average, the model is also better at classifying planets as not having a given molecule when they actually lack it (true negatives). The Perfect Match metric performs poorly, reaching only 50% with SNR greater than 10. Although this result tends to be disappointing, this metric is so stringent that obtaining scores of the order of 0.3 is not as bad as it might seem. Finally, the global Recall seems to reach a maximum value and stops improving significantly when the SNR exceeds 7.
Figure 7: Multilabel classification metrics for the MC as a function of SNR. For the meaning of the shaded region, see Figure 3.
@@ 6.2.2 MC molecule metrics
The novel feature of MC with respect to BC is its ability to allow for a per-molecule classification. In Figure 8, we show the binary metrics for the algorithm when the planets are classified by the presence of each biosignature separately. To calculate the metric, we first choose among the training set those that have a specific value of the SNR and a given molecule, e.g. O3. Then, we use the label, e.g. ozone (that can have two values, 1 or 0 depending on whether the planet has ozone or not, respectively) and the corresponding class assigned by the algorithm, to compute the metrics. Of course, the higher the score the algorithm achieves for a given molecule, the better it is at identifying that molecule among others potentially present in the spectrum.
Figure 8: Molecule-specific metrics for the MC. For the meaning of the shaded region, see Figure 3.
In all metrics, the model shows better performance is attained in classifying O3 compared to CH4 and H2O. Both CH4 and O3 exhibit a recall above 60% even at low SNR. However, as the SNR increases, the Recall remains almost constant. Even though H2O is not a biosignature, it is interesting to notice that the performance of the algorithm in terms of Recall and Precision for this molecule is lower compared to the other molecules. Specifically, H2O generates a trade-off around an SNR of 5, where improving the TNR results in a loss of Recall. This can be interpreted by saying that improving the ability to distinguish planets without water in their atmosphere (i.e. TNR) slightly decreases the ability to detect planets having water.
Comparing all metrics, we notice that the MC algorithm is better at helping us to avoid wasting time with uninteresting candidates, this is because the TNR and Precision metrics have high values (above 80%) for SNR larger than 4.
@@ 6.2.3 MC minimum mixing ratio for detection
Figure 9: Minimum mixing ratio required for detection of each molecule (with Recall $>$ 60%) with MC as a function of SNR values.
As expected, the minimum detectable mixing ratio improves with increasing SNR, even for the case of H2O. However, the minimum mixing ratio for detection of water is larger than for the biosignatures. This is because in the studied spectral range, the absorption strength of water is much lower than that of the other molecules.
Two interesting comparisons between the MC and BC algorithms can be made by examining Figure 4 and Figure 9. While BC is better at detecting low levels of ozone (of the order of $10^{-7}$) at relatively high SNR (greater than 15), in general, MC performs better at lower SNR value (larger or equal than 5). However, MC performs worse because of the interference of the water lines.
@@ 6.2.4 MC interference analysis
Now, we want to evaluate the effect that the presence of one molecule leads to misclassify a planet as if it has another molecule. For instance, at some mixing ratios, the presence of ozone makes the algorithm misclassify a planet as if it has methane although in reality, the latter molecule is not present. This misclassification interference is measured by the TNR metric. We already tested this in subsubsection 6.1.3 for the case of the BC algorithm using water as the interfering molecule.
@@
In Figure 10, we show the result of comparing the TNR measured individually for different molecules (different panels) when the mixing ratio of interference molecules (different curves in each panel) is varied. We show the case for SNR = 3 since, in our numerical experiments, at this value of the SNR the effect of interference was the largest. However, in the numerical repository of the paper (see section 11) we have included animations showing the results for different values of the SNR.
@@
Figure 10: TNR in MC as a function of the mixing ratios for each interfering molecule, at a constant SNR. Each column represents the true negative rate (TNR) for the classification of: CH4 (left), O3 (center), and H2O (right). For the meaning of the shaded region, see Figure 3.
@@
In all cases increasing the mixing ratio of the interference molecule diminished the TNR of each biosignature molecule. In other words, the capability of the algorithm to tag as noninteresting planets that do not have a given molecule is worst when an interference molecule is present. The most significant effect is produced by ozone when classifying a planet by the presence of methane (first panel in Figure 10). At an ozone mixing ratio as low as $10^{-6}$ the algorithm tends to systematically classify planets without methane as if they had it. However, the TNR for O3 (second panel in Figure 10) does not deteriorate as much at higher CH4 concentrations. On the other hand, H2O classification is the most affected by the presence of both biosignatures, showing confusion at various concentrations.
@@
It is interesting to note that the presence of water creates the least interference among all molecules. Only at very high concentrations of water, interference on biosignatures can be significant. This means that this type of algorithm can tag wet planets as inhabited planets, i.e., as if they have biosignatures.
@@
In summary, the MC algorithm faces challenges at classifying molecules individually due to the interference of other molecules. These effects are especially concerning if the mixing ratio of the interfering molecules are relatively high. Interference between ozone and methane is not as concerning since both are biosignatures, however, interference with water, which is not considered a biosignature, can be misleading. Still the mixing ratio of water required for these misclassifications are very large.
@@
A last but not least important experiment we performed with the MC algorithm was to analysing how the presence of other molecules affects the Recall in the classification of a molecule of interest. In Figure 11, we show the Recall maps corresponding to the molecules analysed by our algorithm. This figure is analogous to Figure 6.
@@
Figure 11: Grid of Recall maps for MC at SNR=3. Recall was calculated for different pairs of molecules across varying mixing ratios. Each row represents the classification of a specific molecule (CH4, O3, and H2O from top to bottom), with interference from other molecules as indicated in the labels.
6.3 Specialised Classificator
@@
To address the multilabel classification problem from another perspective, three RF binary classificators were trained, each of them specialised at detecting a specific molecule: $CH_{4}$, $H_{2}O$, or $O_{3}$. The classification of planets according to the presence of all molecules using these individual specialised algorithms will be referred to as Specialised Classificator (SC).
@@
SC is considerably more scalable than previous strategies. On one hand, if we want to add new molecules and/or biosignatures to the classification effort, it only requires training another specialised RF and adding it to the already trained set of algorithms. On the other hand, it demands a lower number and complexity of training spectra as explained below.
@@
Unlike the approach in subsection 6.2, each RF was trained using a dataset containing spectra without biosignatures (fill gases only) and spectra containing fill gases plus the molecule of interest. The numbers of training and test spectra for each molecule depending on the value of SNR are presented in Table 6.
@@
$CH_{4}$ 10000 5000 10000 10000 10000 5000 10000 5000 10000 5000 10000 10000 300000
$O_{3}$ 5000 2500 5000 2500 5000 5000 5000 5000 5000 5000 10000 10000 195000
$H_{2}O$ 10000 5000 10000 5000 10000 5000 10000 5000 10000 5000 10000 10000 285000
Each RF model consisted of 400 estimators that used the entropy criterion, had a maximum depth of 200, and a minimum samples leaf of 3. The classification thresholds for CH4, O3, and H2O were set at 0.40, 0.33, and 0.36, respectively.
To analyse the performance of the SC algorithm at classifying realistic spectra, namely those containing a complex mixture of molecular species (including interfering ones), we generated a spectra dataset using the same procedure described in subsection 6.2. Similarly, the same analyses and plots were generated for the SC as in the case of the MC. This ensures consistency in the assessment of the comparative performance of the models.

### v4

@@ 7.1.3 BC interference analysis
We can use similar methods as before to evaluate the performance of the algorithm in classifying non-bio planets with H2O as interesting cases. This analysis measures the tendency to misclassifying planets due to the presence of water. As shown in Figure 3, the water absorption bands in the spectral range under consideration overlap significantly with most of the ozone bands and some of the methane bands. Consequently, when a planet exhibits a high water mixing ratio and a low SNR, it may be incorrectly classified as interesting.
In Figure 7, we present the TNR (the most appropriate metric for this analysis) as a function of the water mixing ratio. Recall that a high TNR indicates that the algorithm correctly classifies the spectrum of an uninhabited planet as non-interesting. As the figure demonstrates, the TNR decreases as the water mixing ratio increases, which is consistent with the expectation that substantial water interference can lead to misclassification. Conversely, as the SNR increases, the algorithm’s performance improves (i.e. TNR increases), reflecting reduced sensitivity to water interference.
@@
Figure 7: TNR in BC across H2O mixing ratios evaluated at representative SNR values. The shaded area below 0.6 indicates suboptimal classification performance.
@@
We also analyze the performance of the algorithm with planets having two molecules. For that purpose, we picked samples of spectra having pairs of mixing ratios such as (CH4, O3): $(10^{-8},10^{-8})$, $(10^{-8},10^{-7})$, $\ldots$, $(10^{-7},10^{-8})$, $(10^{-7},10^{-7})$, $\ldots$. For each sample, we computed the Recall. In the columns of Figure 8 we show the result as a heat map of this metric as a function of mixing ratio combinations. We call these Recall maps.
@@
Figure 8: Grid of Recall maps for BC at two representative SNR levels (2 and 10). Recall was calculated for different pairs of molecules across varying mixing ratios.
It is important to point out that for these plots we used in the test set spectra with all levels of stellar contamination. That is, while in Figure 7 we classified spectra according to various levels of stellar contamination, for the nature of this 2D heat map we decided to combine all spectra irrespective of the contamination level.
@@
We observe that even at low mixing ratios of one of the biosignatures/bioindicators, Recall can be high if the mixing ratio of the other is high enough. For example, at SNR = 2 (first row in Figure 8) and at a very low concentration of ozone (e.g. $10^{-8}$ in the first panel), the presence of abundant methane (e.g. $10^{-3}$) causes the algorithm to classify as interesting most planets that exhibit both biosignatures/bioindicators. Our interpretation is that the algorithm learns the spectral bands of each biosignature/bioindicator independently and is able to recognise that a planet is interesting by identifying any one of them, even if the other is absent.
@@
Additionally, as SNR increases, the algorithm becomes more conservative (yielding lower Recall values) when both mixing ratios are low, but achieves better performance (higher Recall values) at higher mixing ratios.
@@
We repeated the same procedure using combinations of one biosignature/bioindicator and water. The resulting Recall maps are shown in the second and third columns of Figure 8. Although water tends to interfere with the detection of biosignatures/bioindicators, the fact that the frontier between low and high Recall areas in the Recall maps is nearly horizontal indicates that the performance of the algorithm is mainly determined by the mixing ratio of the biosignature. In other words, H2O only negatively affects the Recall when the mixing ratio of the biosignature is very low.
7.1.4 BC for airless and CO2 rich planets
To evaluate the robustness of the binary classifier against false positives, its performance was investigated in two critical scenarios: (i) airless planets, where noise and contamination may lead to the misinterpretation of spectral features as molecular bands when no atmosphere is actually present; and (ii) atmospheres composed of fill gases, such as CO2 and N2, for which spectral bands fall in the same region as the sought bioindicator—particularly in the case of O3—and this “virtual overlapping” could result in erroneous classifications.
In the airless scenario, the TNR was always larger than 0.9 even at an SNR of 1 and under the most adverse stellar contamination. This means that if we take 100 test spectra of airless planets, more than 90 will be correctly classified as non-interesting. By contrast, for pure fill gas atmospheres, the TNR is approximately 0.8 at an SNR of 1, improving to values exceeding 0.9 at an SNR of 3.
These findings demonstrate that the binary classifier is robust at avoiding false positives.
7.2 Multilabel Classifier
@@
In our previous experiment we binary classified the planets as interesting and not-interesting. We now aim to improve the method by introducing multiple labels intended to identify individual biosignatures/bioindicators and other interesting molecules.
@@
For that purpose, we specifically define three labels: methane, ozone and water, each with two possible values: 0 (the molecule is absent) and 1 (the molecule is present). Focusing on the detection of each molecule individually is valuable because, as discussed in section 4, different biosignatures/bioindicators have varying degrees of robustness and implications. By independently classifying each biosignature, we can better understand the presence and relevance of each molecule, thereby improving our ability to assess the potential habitability of exoplanets.
@@
The MC random forest uses the same hyperparameters as the BC algorithm described in subsection 7.1 (100 estimators and a maximum depth of 200).
@@
Table 5: Training data distribution for the multilabel classifier (MC). This table has the same structure as Table 4, with each cell’s value (in thousands of samples) representing the total number of noisy spectra generated for that atmospheric configuration.
SNR 000 100 010 001 110 101 011 111 airless
1 300 60 60 60 57.6 57.6 57.6 76.8 100
3 300 60 60 60 57.6 57.6 57.6 76.8 100
6 300 24 24 24 38.4 38.4 38.4 46.08 100
10 300 24 24 24 38.4 38.4 38.4 46.08 100
0 300 60 60 60 96 96 96 76.8 100
After training, the classification thresholds for CH4, O3, and H2O were set at 0.28, 0.24, and 0.36, respectively.
The model was tested by generating approximately 750,000 noisy spectra per SNR. This included 150,000 spectra for the scenario without interesting molecules (comprising only filler gases and airless cases, in equal proportions), and another 150,000 spectra for the cases where all three molecules are present. For the remaining combinations, we generated 75,000 spectra per combination, resulting in a total of 450,000 spectra. This distribution ensures a balanced test dataset, meaning that for each molecule there is an equal number of spectra where it is present and absent.
@@ 7.2.1 MC multilabel metrics
Figure 9: Multilabel classification metrics for the MC as a function of SNR. For the meaning of the shaded regions, see Figure 5.
As we see, the overall performance of the algorithm is very good for most metrics even at an SNR of 1, and it improves further as the SNR increases. The only metric that performs poorly is the Perfect Match, which measures the algorithm’s capacity to correctly classify the composition of a planet’s atmosphere with the perfect label—irrespective of the combination of molecular species (see Table 2). For the MC algorithm, the Perfect Match does not exceed 50% and is only achieved when the SNR is 10 or higher.
Although this result may seem disappointing, it is important to note that this metric is extremely stringent; thus, scores on the order of 0.3 are not as adverse as they might appear.
In this figure, we have also included a special Recall metric, termed Chemical Disequilibrium Recall, which measures the algorithm’s capacity to label a planet as having both CH4 and O3 simultaneously. This is an interesting case since the presence of both gases indicates a chemical disequilibrium—a condition recognised as one of the strongest evidences for an inhabited planet (see section 4). Notably, the MC algorithm is able to flag more than 70% of the inhabited planets in our test set as interesting cases for follow-up observations, even at an SNR of 1.
To illustrate this point about how impressive this achievement is, please refer to Figure 4, which shows examples of the types of spectra we are dealing with. It is particularly noteworthy that the MC algorithm achieves accurate classification of most spectra, even under conditions of very low spectral feature contrast, which few retrieval algorithms would reliably fit.
@@ 7.2.2 MC molecule metrics
The novel feature of MC with respect to BC is its ability to allow for per-molecule classification. In Figure 10, we show the binary metrics for the algorithm when the planets are classified by the presence of each bioindicator separately. To calculate the metric, we first select, from the training set, the spectra with a specific SNR value and containing a given molecule (e.g. O3). Then, using the corresponding label (e.g. ozone, which can have a value of 1 if the planet has ozone or 0 otherwise) and the class assigned by the algorithm, we compute the metrics. Naturally, the higher the score achieved for a given molecule, the better the algorithm is at identifying that molecule among others that may be present in the spectrum.
Figure 10: Molecule-specific metrics for the MC. The shaded regions indicate the variance in performance (see Figure 5 for details).
In all metrics, better performance is attained when classifying O3 (red continuous line and shaded strip) compared to CH4 and H2O (blue strip). Both CH4 and O3 exhibit a Recall above 60% even at low SNR.
Comparing all metrics, we observe that the MC algorithm is effective at identifying interesting candidates. This is evident from the Recall (first row, second column in Figure 10), which reaches values above 80% for SNRs greater than 3. Furthermore, the TNR and Precision metrics—which help avoid wasting time on uninteresting targets—are also satisfactory. In fact, for the interesting bioindicators, even at low SNR values of 3–4, these metrics are close to or exceed 0.8, depending on the level of stellar contamination (as indicated by the shaded areas around the continuous lines).
@@ 7.2.3 MC minimum mixing ratio for detection
Figure 11: Minimum mixing ratio required for detection of each molecule (with Recall > 60%) by MC as a function of SNR.
As expected, the minimum detectable mixing ratio decreases with increasing SNR, even for H2O. However, the minimum mixing ratio required for the detection of water is higher than that for the biosignatures/bioindicators. This is because, in the studied spectral range, the absorption strength of water is much lower than that of the other molecules.
@@ 7.2.4 MC interference analysis
Now, we aim to evaluate the effect of one molecule’s presence on causing the misclassification of a planet as if it contains another molecule. For instance, at certain mixing ratios, the presence of ozone causes the algorithm to misclassifying a planet as if it contains methane, even though methane is absent. This misclassification interference is measured by the TNR metric. We already tested this phenomenon in subsubsection 7.1.3 for the BC algorithm using water as the interfering molecule.
@@
In Figure 12, we present the results of comparing the TNR measured individually for different molecules (in separate panels) while varying the mixing ratio of the interfering molecules (represented by different curves in each panel). We focus on the case for SNR = 3, as our numerical experiments indicate that the interference effect is most pronounced at this SNR value.
@@
Figure 12: TNR in MC as a function of the mixing ratios for each interfering molecule, at a constant SNR. Each column represents the true negative rate (TNR) for the classification of: CH4 (left), O3 (center), and H2O (right). For the meaning of the shaded regions, see Figure 5.
@@
In all cases, increasing the mixing ratio of the interfering molecule reduced the TNR for each biosignature/bioindicator. In other words, the algorithm’s ability to correctly label as non-interesting those planets that lack a given molecule is poorest when an interfering molecule is present. The most significant effect is observed when ozone interferes with the classification of methane (first panel in Figure 12). At an ozone mixing ratio as low as $10^{-6}$$10^{-5}$, the algorithm tends to systematically classify planets that do not contain methane as if they did. However, the TNR for O3 (second panel in Figure 12) does not decline as much at higher CH4 concentrations. In contrast, the classification of H2O is most affected by the presence of both biosignatures/bioindicators, leading to confusion across a range of concentrations.
@@
It is interesting to note that the presence of water creates the least interference among all molecules. Only at very high water concentrations (exceeding approximately $10^{-4}$) does interference with biosignatures/bioindicators become significant. This implies that the algorithm may erroneously tag wet planets as inhabited—i.e. as possessing biosignatures/bioindicators—even if they do not.
@@
In summary, the MC algorithm faces challenges in classifying molecules individually due to interference from other molecules. These effects are particularly concerning when the mixing ratios of the interfering molecules are relatively high. Interference between ozone and methane is less concerning, since both are considered biosignatures/bioindicators; however, interference involving water—which is not regarded as a biosignature—can be misleading. Nonetheless, the water mixing ratios required for such misclassifications are relatively large [^3] For present-day Earth, the integrated water mixing ratio is around $10^{-5}$ (Lustig-Yaeger et al., 2023), depending on the atmospheric model.
@@
A final, but equally important, experiment we performed with the MC algorithm was to analyse how the presence of other molecules affects the Recall in the classification of a molecule of interest. In Figure 13, we present the Recall maps for the molecules analysed by our algorithm. This figure is analogous to Figure 8; however, in contrast to the latter, a single SNR value of 3 was used to illustrate the interference effect.
@@
Figure 13: Grid of Recall maps for MC at SNR=3. Recall was calculated for different pairs of molecules across varying mixing ratios. Each row represents the classification for a specific molecule (CH4, O3, and H2O, from top to bottom), with interference from other molecules as indicated in the labels.
In all cases, contrary to what might be expected, interference may actually enhance the detection of a given molecule. This is reflected in the figure as darker areas in the lower-left corner of each panel. This effect is particularly pronounced for methane; even at very low methane mixing ratios, the presence of high concentrations of ozone can cause the algorithm to label a planet as having methane (as seen in the lower-right clear region of the corresponding panel). This behaviour is consistent with the TNR results shown in Figure 12. An ideal Recall map would display a dark strip at low mixing ratios of the target molecule; this is nearly the case for ozone when analysed for interference from water (second panel in the second row).
7.2.5 MC for airless and CO2-rich planets
Following a similar analysis to that described in subsubsection 7.1.4, but now applied to the classification of individual molecules (CH4, O3, and H2O), we evaluated the performance of MC in distinguishing two background scenarios: airless planets and CO2-dominated atmospheres without other molecules. For airless scenarios, TNR values exceed 0.8 even in the most contamination-affected classifications at an SNR of 1, although significant improvement is only observed at an SNR of 5, where TNR values approach 0.9 for all molecules, even in the most challenging cases. In the case of CO2-rich atmospheres, the classifier also exhibits robust performance. At an SNR of 1, TNR values remain above 0.7 in the worst-case scenario, with CH4 and H2O achieving values above 0.75, while O3 drops to 0.7—likely due to the overlap between its main absorption peak and one of the CO2 spectral bands. However, at an SNR of 3, TNR values exceed 0.8 for all molecules, confirming that the model maintains reliable classification performance under stellar contamination and in CO2-dominated atmospheric conditions.
7.3 Specialised Classifier
@@
To address the multilabel classification problem from another perspective, three RF binary classifiers were trained, each specialised in detecting a specific molecule: CH4, H2O, or O3. The classification of planets according to the presence of all molecules using these individual specialised algorithms will be referred to as the Specialised Classifier (SC).
@@
SC is considerably more scalable than previous strategies. On one hand, if we want to add new molecules and/or bioindicators to the classification effort, it only requires training another specialised RF and adding it to the existing set of algorithms. On the other hand, it demands a lower number and complexity of training spectra, as explained below.
@@
Unlike the approach in subsection 7.2, each RF was trained using a dataset containing spectra of airless planets, planets without bioindicators (i.e., only filler gases), and planets containing filler gases plus the molecule of interest. The numbers of training and test spectra for each molecule, depending on the SNR value, are presented in Table 6.
@@
Molecule 400000 240000 280000 240000 200000 120000 200000 120000 480000 400000 2680000
Each RF model consisted of 100 estimators that used the entropy criterion, had a maximum depth of 200, and a minimum samples leaf of 3. The classification thresholds for CH4, O3, and H2O were set at 0.32, 0.24, and 0.37, respectively (for an explanation about the selection of the classification thresholds, see section 7).
To analyse the performance of the SC algorithm in classifying realistic spectra—namely, those containing a complex mixture of molecular species (including interfering ones)—we generated a test dataset of spectra using the same procedure described in subsection 7.2. Similarly, the same analyses and plots were generated for the SC as in the case of the MC. This ensures consistency in the comparative assessment of the models’ performance.